# Audio-Visual Instance Segmentation

### v2

@@ Abstract
In this paper, we propose a new multi-modal task, termed audio-visual instance segmentation (AVIS), which aims to simultaneously identify, segment and track individual sounding object instances in audible videos. To facilitate this research, we introduce a high-quality benchmark named AVISeg, containing over 90K instance masks from 26 semantic categories in 926 long videos. Additionally, we propose a strong baseline model for this task. Our model first localizes sound source within each frame, and condenses object-specific contexts into concise tokens. Then it builds long-range audio-visual dependencies between these tokens using window-based attention, and tracks sounding objects among the entire video sequences. Extensive experiments reveal that our method performs best on AVISeg, surpassing the existing methods from related tasks. We further conduct the evaluation on several multi-modal large models; however, they exhibits subpar performance on instance-level sound source localization and temporal perception. We expect that AVIS will inspire the community towards a more comprehensive multi-modal understanding.
@@ 1 Introduction
Vision and hearing are our primary channels of communication and sensation. Audio-visual collaboration is beneficial for humans to better perceive and interpret the world. Humans have the ability to associate mixed sounds with object instances in complicated realistic scenarios. Imagine a cocktail-party scenario: when a group of people is speaking, we can not only locate the sound sources but also determine how many people are talking.
Inspired by this human perception, we explore instance-level sound source localization in long videos and propose a new task, namely audio-visual instance segmentation (AVIS). As can be seen in Figure 1 (c), it requests a model to simultaneously classify, segment and track sounding object instances—identify which objects are making sounds, infer where the sounding objects are, and monitor when they are making sounds. This new task facilitates a wide range of practical applications, including embodied robotics, virtual reality, video surveillance, video editing, etc. Moreover, it can serve as a fundamental task for evaluating the comprehension capabilities of multi-modal large models.
@@
To deal with the above AVIS task, we follow the query-based segmentation paradigm [9, 22] and propose a baseline model called AVISM. To be specific, a frame-level sound source localizer segments sounding objects within each frame independently and summarizes per-frame scenes into a small amount of object tokens. Then, a video-level sounding object tracker is designed to build frame-to-frame communications and track sounding objects throughout the entire video. To lessen computational overheads in processing long and high-resolution videos, the tracker uses the concise object tokens as a mean of conveying information rather than dense image features, and adopts window-based self-attention mechanisms to efficiently capture long-range dependencies in consecutive frames. Experimental results demonstrate the superiority of our baseline model. Additionally, we make a thorough evaluation of several prominent multi-modal large models on our AVISeg dataset. Surprisingly, these self-proclaimed large models are far from satisfactory in instance-aware sound source localization and temporal perception. Our dataset emphasizes the necessity for further improvements in handling audio-visual data and long videos, providing valuable insights for future development of multi-modal large models.
To summarize, our contributions are fourfold:
@@ 2.2 Audio-Visual Segmentation
Audio-visual segmentation (AVS) focuses on localizing and segmenting sounding objects within each video frame. Zhou et al. [57, 58] introduce the first AVS dataset, namely AVSBench, which serves two different sub-tasks including audio-visual object segmentation (AVOS) and audio-visual semantic segmentation (AVSS). The former [57] requires producing binary masks of sounding objects, while the latter [58] further needs to generate semantic maps representing the object category. To address these problems, they employ a standard encoder-decoder architecture with a modified non-local block to encode space-time relation and segment sounding objects. ECMVAE [37] factorizes the representations of audio and visual data and maximizes the contribution of each modality for the AVOS task. Inspired by the success of DETR [5] and Mask2Former [9], recent works [14, 50, 31, 19, 18] adopt the query-based architecture decode masks for sounding objects. For example, AVSegFormer [14] trivially incorporates audio features and learnable queries, enabling the decoder to capture relevant visual semantics and predict the audio-constrained masks. COMBO [50] explores multi-order bilateral relations in modality, temporal and pixel levels for the AVSS task. Notably, a bilateral-fusion module is designed to align audio and visual modalities bi-directionally and assist the model in segmenting the sounding objects.
@@ 3.2 Evaluation Metrics
In order to evaluate how well an AVIS model performs, we need to choose appropriate metrics to compare its outputs with the ground truth. In the era of competitive benchmarks, much research is evaluated on its ability to improve the scores on these benchmarks. If benchmarks are using metrics to evaluate these scores which are skewed towards only certain aspects of a task, this will also steer research and models to focus more on these aspects. In our task, we adopt two evaluation protocols including the mean Average Precision (mAP) [49] and the Higher Order Tracking Accuracy (HOTA) [36]. mAP follows the computation of the average precision-recall metric over trajectories, which is commonly used in video instance segmentation. However, mAP is not perfectly suited to our task, because it can be increased by producing many different predictions with low confidence scores and does not decrease even if non-sounding objects are predicted. HOTA performs a bijective (one-to-one) matching at the detection level while scoring association globally over trajectories, which is designed for multi-object tracking task. This makes HOTA a balanced metric for measuring both detection and association. When applied to the AVIS task, it can penalize those models that predict non-sounding objects.
@@ 4 Dataset
To explore audio-visual instance segmentation and evaluate the proposed methods, we create a new large-scale benchmark called AVISeg. Considering that this task involves complex audio-visual interactions and requires high-quality data, we manually collect and choose 926 videos from YouTube and the publicly available MUSIC-AVQA [29] dataset. Our released AVISeg dataset satisfies the following criteria: 1) It focuses on long-term videos, bringing them much closer to real applications. 2) It contains 26 common sound categories, spanning 4 dynamic scenarios: “Music”, “Speaking”, “Machine”, and “Animal”. 3) It involves some challenging cases, such as videos with silent sound sources, single sound source, and multiple sources simultaneously. These attributes impose higher demands on the model for accurate recognition, segmentation, and tracking of sounding objects.
@@
Table 1: Comparison with other related datasets. SSL represents audio-visual event localization.
Task Dataset Videos
Video
Length
Classes Anno.
@@ 6.3 Ablation Studies
Impact of visual backbone and pre-training dataset. We further investigate whether providing a stronger backbone and more pre-training data can further improve the model’s AVIS performance. As shown in Table 6, adopting the strategy from Mask2Former [9] that using COCO for additional pre-training of our visual backbone resulted in improvements across all metrics. However, when further fine-tuned on the video instance segmentation dataset OVIS [39], despite an increase in mAP, we observe a slight decrease in FSLA. This is likely because OVIS primarily targets improving the model’s video segmentation capabilities, leading to the segmentation of many non-sounding objects, thus not achieving better FSLA scores. Consequently, we opt for the IN+COCO pre-trained visual backbone for subsequent experiments. Replacing the backbone with Swin-L further enhances AVIS performance. Specifically, the Swin-L backbone pre-trained on IN+COCO achieves the highest scores across all metrics.
@@ 7 Conclusion
8 Overview
9 Evaluation Metrics
10 Implementation Details
11 Details of Multi-modal Large Models
12 More Qualitative Results
13 Failure Cases
14 Future Works
9 Evaluation Metrics
9.1 mAP: mean Average Precision
@@ 9 Evaluation Metrics
9.2 HOTA: Higher Order Tracking Accuracy
@@
9.3 FSLA: Frame-level Sound Location Accuracy
@@
Input : A predicted mask set $\mathbf{M^{P}}=\{\mathbf{m^{P}_{i}}\}_{i=1}^{x}$.
A labeled mask set $\mathbf{M^{G}}=\{\mathbf{m^{G}_{j}}\}_{j=1}^{y}$.
@@
4 $S(x,y)=HungarianMatch(\mathbf{M^{P}},\mathbf{M^{G}},\mathbf{ID^{P}},\mathbf{ID^{G}})$ for $\alpha\leftarrow 0.05$ to $0.95$ step $0.05$ do
@@ 7                   $N_{fnt}\leftarrow N_{fnt}+1$
10                   if $IoU(\mathbf{m^{P}_{1}},\mathbf{m^{G}_{1}})>\alpha$ then
@@ 11                         $N_{fst}\leftarrow N_{fst}+1$
16                   if $IoU(\mathbf{m^{P}_{i}},\mathbf{m^{G}_{j}})>\alpha$ and $S(i,j)$ then
@@ 17                         $N_{fmt}\leftarrow N_{fmt}+1$
23      FSLAn($\alpha$) $\leftarrow$ $N_{fnt}/N_{fna}$
24       FSLAs($\alpha$) $\leftarrow$ $N_{fst}/N_{fsa}$
@@ 26       FSLA($\alpha$) $\leftarrow$ $(N_{fnt}+N_{fst}+N_{fmt})/L$
29FSLA $\leftarrow$ $\overline{\text{FSLA}(\alpha)}$
@@ 10 Implementation Details
11 Details of Multi-modal Large Models
11.1 Sam4AVS
@@ 11.4 AL-Ref-SAM 2
To avoid the disturbance presence of background noise and the ambiguity of audio information, it integrates visual context and leverages the vision-language understanding capabilities of existing large multi-modal model, GPT-4 [1], to accurately identify the categories of the actual sounding objects present in the video. The prompt template, i.e., prompt1 depicted in Figure 6 (d), is defined as mentioned above. Since the selected referent may be silent in certain frames, AL-Ref-SAM 2 further utilizes sound event detection (SED) to segment the whole audio clip and filter out silent frames from the generated mask sequence. Then, GPT-4 processes the identified categories and video clip to identify a high-quality box of the referent in a specific frame where the referent clearly appears. The prompt template, i.e., prompt2 depicted in Figure 6 (d), is defined as follows:
@@ 4. Identify the Object that Best Matches the Description:
Finally, the selected bounding box serves as the pivot box to effectively prompt SAM 2 [41] to segment the referent and propagate its mask forward and backward through the entire video clip.
@@
12 More Qualitative Results
As shown in Figure 7 and Figure 8, we provide some qualitative comparisons with other methods on 4 scenarios. 1) Video instance segmentation methods (e.g., VITA [22]) can accurately segment and track objects, but fails to determine when these objects are producing sound, e.g., “person” in Figure 7 and “lion” in Figure 8, due to the absence of audio input. 2) With the help of audio information, audio-visual semantic segmentation methods (e.g., COMBO [50]) are capable of correctly localizing the sound source in most cases. However, such methods show difficulties in processing long sequences, which may result in multiple identity switches in tracking, e.g., “person” in Figure 7. 3) Multi-modal large models (e.g., PG-Video-LLaVA [38] and AL-Ref-SAM 2 [23]) serve audio as a form of language and leverage foundation models to achieve audio-referred visual grounding. As discussed in Section 11, these methods not only fail to distinguish between different object instances of the same category, e.g., “person” in Figure 7, but also struggle to determine the exact time intervals for each sounding object, e.g., “lion” in Figure 8.
@@ 13 Failure Cases
14 Future Works
2) Audio decoupling and audio-visual fusion. As discussed in Section 13, our model’s performance may be limited in scenarios where multiple objects of the same category are producing sound. To better associate mixed-source audio with visual objects, product quantization [25, 50] can be considered to decompose the mixed audio semantics into several disentangled single-source semantics with noise suppression. This approach has the potential to provide a more compact and robust audio representation for audio-visual interaction, especially in complex scenarios.

### v4

@@ Abstract
In this paper, we propose a new multi-modal task, termed audio-visual instance segmentation (AVIS), which aims to simultaneously identify, segment and track individual sounding object instances in audible videos. To facilitate this research, we introduce a high-quality benchmark named AVISeg, containing over 90K instance masks from 26 semantic categories in 926 long videos. Additionally, we propose a strong baseline model for this task. Our model first localizes sound source within each frame, and condenses object-specific contexts into concise tokens. Then it builds long-range audio-visual dependencies between these tokens using window-based attention, and tracks sounding objects among the entire video sequences. Extensive experiments reveal that our method performs best on AVISeg, surpassing the existing methods from related tasks. We further conduct the evaluation on several multi-modal large models. Unfortunately, they exhibits subpar performance on instance-level sound source localization and temporal perception. We expect that AVIS will inspire the community towards a more comprehensive multi-modal understanding. Dataset and code is available at https://github.com/ruohaoguo/avis.
@@ 1 Introduction
Vision and hearing are our primary channels of communication and sensation [66, 68, 67, 64, 63, 9, 23]. Audio-visual collaboration is beneficial for humans to better perceive and interpret the world. Humans have the ability to associate mixed sounds with object instances in complicated realistic scenarios. Imagine a cocktail-party scenario: when a group of people is speaking, we can not only locate the sound sources but also determine how many people are talking.
Inspired by this human perception, we explore instance-level sound source localization in long videos and propose a new task, namely audio-visual instance segmentation (AVIS). As can be seen in Figure 1 (c), it requests a model to simultaneously classify, segment and track sounding object instances—identify which object categories are making sounds, infer where the sounding objects are, and monitor when they are making sounds. This new task facilitates a wide range of practical applications, including embodied robotics, video surveillance, video editing, etc. Moreover, it can serve as a fundamental task for evaluating the comprehension capabilities of multi-modal large models.
@@
In order to deal with the above AVIS task, we follow the query-based segmentation paradigm [12, 26] and propose a baseline model called AVISM. To be specific, a frame-level sound source localizer segments sounding objects within each frame independently and summarizes per-frame scenes into a small amount of object tokens. Then, a video-level sounding object tracker is designed to build frame-to-frame communications and track sounding objects throughout the entire video. To lessen computational overheads in processing long and high-resolution videos, the tracker uses the concise object tokens as a mean of conveying information rather than dense image features, and adopts window-based self-attention mechanisms to efficiently capture long-range dependencies in consecutive frames. Experimental results demonstrate the superiority of our baseline. Additionally, we make a thorough evaluation of several prominent multi-modal large models on our AVISeg dataset. Surprisingly, these self-proclaimed large models are far from satisfactory in instance-aware sound source localization and temporal perception. Our dataset emphasizes the necessity for further improvements in handling audio-visual data and long videos, providing insights for future development of multi-modal large models. Our contributions are as follows:
@@ 2.2 Audio-Visual Segmentation
Audio-visual segmentation (AVS) focuses on localizing and segmenting sounding objects within each video frame. Zhou et al. [65, 69] introduce the first AVS dataset, namely AVSBench, which serves two different sub-tasks including audio-visual object segmentation (AVOS) and audio-visual semantic segmentation (AVSS). The former [65] requires producing binary masks of sounding objects, while the latter [69] further needs to generate semantic maps representing the object category. To address these problems, they employ a standard encoder-decoder architecture with a modified non-local block to encode space-time relation and segment sounding objects. CAVP [10] builds an AVS dataset by randomly matching the images from COCO [38] and audio files from VGGSound [6] based on the semantic classes of the objects. Inspired by DETR [5] and Mask2Former [12], recent works [17, 56, 35, 22, 21] adopt the query-based architecture decode masks for sounding objects. For example, AVSegFormer [17] trivially incorporates audio features and learnable queries, enabling the decoder to capture relevant visual semantics and predict the audio-constrained masks. COMBO [56] explores multi-order bilateral relations in modality, temporal and pixel levels for the AVSS task. Notably, a bilateral-fusion module is designed to align audio and visual modalities bi-directionally and assist the model in segmenting the sounding objects.
@@ 3.2 Evaluation Metrics
To evaluate how well an AVIS model performs, we need to choose appropriate metrics to compare its outputs with the ground truth. In our task, we adopt two evaluation protocols including the mean Average Precision (mAP) [55] and the Higher Order Tracking Accuracy (HOTA) [43]. mAP follows the computation of the average precision-recall metric over trajectories, which is commonly used in video instance segmentation. However, mAP is not perfectly suited to our task, because it can be increased by producing many different predictions with low confidence scores and does not decrease even if non-sounding objects are predicted. HOTA performs a bijective matching at the detection level while scoring association over trajectories, which is designed for multi-object tracking task. This makes HOTA a balanced metric for measuring both detection and association. When applied to the AVIS task, it can penalize those models that predict non-sounding objects.
@@ 4 Dataset
To explore audio-visual instance segmentation and evaluate the proposed methods, we create a new large-scale benchmark called AVISeg. Considering that this task involves complex audio-visual interactions and requires high-quality data, we manually collect and choose 926 videos from YouTube and the publicly available datasets [33, 36, 37], e.g., MUSIC-AVQA. Our released AVISeg dataset satisfies the following criteria: 1) It focuses on long-term videos (61.4s), bringing them much closer to real applications. 2) It contains 26 common sound categories, spanning 4 dynamic scenarios: “Music”, “Speaking”, “Machine”, and “Animal”. 3) It involves some challenging cases, such as videos with silent sound sources, single sound source, and multiple sources simultaneously. These attributes impose higher demands on the model for accurate recognition, segmentation, and tracking of sounding objects.
@@
Table 1: Comparison with other datasets from related tasks. SSL represents audio-visual event localization.
Task Dataset Videos Length Classes Anno.
@@ 6.3 Ablation Studies
Impact of visual backbone and pre-training dataset. We further investigate whether providing a stronger backbone and more pre-training data can further improve the model’s AVIS performance. As shown in Table 6, adopting the strategy from Mask2Former [12] that using COCO for additional pre-training of our visual backbone resulted in improvements across all metrics. However, when further fine-tuned on the video instance segmentation dataset OVIS [45], despite an increase in mAP, we observe a slight decrease in FSLA. This is likely because OVIS primarily targets improving the model’s video segmentation capabilities, leading to the segmentation of many non-sounding objects, thus not achieving better FSLA scores. Consequently, we opt for the IN+COCO pre-trained visual backbone for subsequent experiments. Replacing the backbone with Swin-L achieves the highest scores across all metrics.
@@ 7 Conclusion
Acknowledgments. This work was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62371009, and Beijing Natural Science Foundation under Grant No. L247029.
@@
Appendix A Overview
B Evaluation Metrics
C Implementation Details
D More Ablation Studies
E Details of Multi-modal Large Models
F More Qualitative Results
G Failure Cases
H Future Works
Appendix B Evaluation Metrics
In the era of competitive benchmarks, much research is evaluated on its ability to improve the scores. If benchmarks are using metrics to evaluate these scores which are skewed towards only certain aspects of a task, this will also steer research and models to focus more on these aspects.
B.1 mAP: mean Average Precision
@@ Appendix B Evaluation Metrics
B.2 HOTA: Higher Order Tracking Accuracy
@@
B.3 FSLA: Frame-level Sound Location Accuracy
@@
Input : A predicted mask set $\mathbf{M^{P}}=\{\mathbf{m^{P}_{i,l}}\}_{i=1,l=1}^{x,L}$.
A labeled mask set $\mathbf{M^{G}}=\{\mathbf{m^{G}_{j,l}}\}_{j=1,l=1}^{y,L}$.
@@
4 $S(x,y)=\text{{HungarianMatch}}(\mathbf{M^{P}},\mathbf{M^{G}},\mathbf{ID^{P}},\mathbf{ID^{G}})$
5 for $\alpha\leftarrow 0.05$ to $0.95$ step $0.05$ do
@@ 8                   $N_{fnt}\leftarrow N_{fnt}+1$
11                   if $\mathbf{c^{P}_{i,l}}==\mathbf{c^{G}_{j,l}}$ then
12                         if $S(i,j)$ and $IoU(\mathbf{m^{P}_{i,l}},\mathbf{m^{G}_{j,l}})>\alpha$ then
@@ 13                               $N_{fst}\leftarrow N_{fst}+1$
20                   if $\mathbf{c^{P}_{i,l}}==\mathbf{c^{G}_{j,l}}$ then
21                         if $S(i,j)$ and $IoU(\mathbf{m^{P}_{i,l}},\mathbf{m^{G}_{j,l}})>\alpha$ then
@@ 22                               $N_{fmt}\leftarrow N_{fmt}+1$
30      FSLAn($\alpha$) $\leftarrow$ $N_{fnt}/N_{fna}$, FSLAs($\alpha$) $\leftarrow$ $N_{fst}/N_{fsa}$
@@ 32       FSLA($\alpha$) $\leftarrow$ $(N_{fnt}+N_{fst}+N_{fmt})/L$
35FSLA $\leftarrow$ $\overline{\text{FSLA}(\alpha)}$
@@ Appendix C Implementation Details
Appendix D More Ablation Studies
Impact of similarity loss and hyper-parameter setup. As shown in Table 7, removing similarity loss yields a significant decrease across all metrics. This is because the model struggles to learn correct associations between object tokens and video queries, leading to feature misalignment, identity switches and tracking failures, especially for different instances of the same category. Additionally, we test several hyper-parameters and set $\lambda_{\text{sim}}=0.5$ as default, which achieves the best performance.
Table 7: Impact of similarity loss and hyper-parameter setup.
similarity loss $\lambda_{\text{sim}}=0.1$ $\lambda_{\text{sim}}=0.5$ $\lambda_{\text{sim}}=1.0$ FSLA HOTA mAP
[…]
Appendix E Details of Multi-modal Large Models
E.1 Sam4AVS
@@ 4. Identify the Object that Best Matches the Description:
To avoid the disturbance presence of background noise and the ambiguity of audio information, it integrates visual context and leverages the vision-language understanding capabilities of existing large multi-modal model, GPT-4 [1], to accurately identify the categories of the actual sounding objects present in the video. The prompt template, i.e., prompt1 depicted in Figure 6 (d), is defined as mentioned above. Since the selected referent may be silent in certain frames, AL-Ref-SAM 2 further utilizes sound event detection (SED) to segment the whole audio clip and filter out silent frames from the generated mask sequence. Then, GPT-4 processes the identified categories and video clip to identify a high-quality box of the referent in a specific frame where the referent clearly appears. The prompt template, i.e., prompt2 depicted in Figure 6 (d), is defined as above. Finally, the selected bounding box serves as the pivot box to prompt SAM 2 [47] to segment the referent and propagate its mask forward and backward through the entire video.
@@
Appendix F More Qualitative Results
As shown in Figure 7 and Figure 8, we provide some qualitative comparisons with other methods on 4 scenarios. 1) Video instance segmentation methods (e.g., VITA [26]) can accurately segment and track objects, but fails to determine when these objects are producing sound, e.g., “person” in Figure 7 and “lion” in Figure 8, due to the absence of audio input. 2) With the help of audio information, audio-visual semantic segmentation methods (e.g., COMBO [56]) are capable of correctly localizing the sound source in most cases. However, such methods show difficulties in processing long sequences, which may result in multiple identity switches in tracking, e.g., “person” in Figure 7. 3) Multi-modal large models (e.g., PG-Video-LLaVA [44] and AL-Ref-SAM 2 [27]) serve audio as a form of language and leverage foundation models to achieve audio-referred visual grounding. As discussed in Section E, these methods not only fail to distinguish between different object instances of the same category, e.g., “person” in Figure 7, but also struggle to determine the exact time intervals for each sounding object, e.g., “lion” in Figure 8.
@@ Appendix F More Qualitative Results
Appendix G Failure Cases
Appendix H Future Works
2) Audio decoupling and audio-visual fusion. As discussed in Section G, our model’s performance may be limited in scenarios where multiple objects of the same category are producing sound. To better associate mixed-source audio with visual objects, product quantization [29, 56] can be considered to decompose the mixed audio semantics into several disentangled single-source semantics with noise suppression. This approach has the potential to provide a more compact and robust audio representation for audio-visual interaction, especially in complex scenarios.